EMPLOYING DIFFUSION MONTE CARLO IN THE CALCULATION OF MINIMIZED ENERGY PATHS OF THE CH_{3}^{+} + H_{2} ↔ CH_{5}^{+} ↔ CH_{3}^{+} + H_{2} REACTION AND ITS ISOTOPIC VARIANTS

\underline{CHARLOTTE E. HINKLE}, ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Protonated methane is presumed by astrochemists to be an important intermediate in the reaction CH_{3}^{+} + HD → CDH_{4}^{+} → CH_{2}D^{+} + H_{2} within the interstellar medium. Understanding this reaction can also help shed light on the observed nonstatistical H/D isotopic abundance in the isotopologues of CH_{3}^{+} within the interstellar medium. Interestingly, based on kinetic studies, Gerlich and co-workers showed that all of the reactions in the series CH_{3−n}D_{n}^{+} + HD → CH_{4−n}D_{n+1}^{+} → CH_{2−n}D_{n+1}^{+} + H_{2} have identical net rate constants .^{a} This result is independent of the value of n.

In previous studies of CH_{5}^{+}, we have employed Diffusion Monte Carlo (DMC)^{b} to study ground,^{c} and excited states.^{d}^{,}^{e}^{,}^{f} By performing the simulation in Jacobi coordinates, we can use Adiabatic DMC^{g} to study the properties of the minimized energy paths of CH_{5}^{+} and isotopologues. To determine the minimized energy path, we calculate the quantum zero-point energy and ground state wave function as a function of the distance between the center of mass of the H_{2} group and the center of mass of the CH_{3}^{+} group over a range from 0 to 6 Å. Over this range, we find 5 distinct regions of interaction, short range repulsion region, CH_{5}^{+} complexation, short-range fragment interaction, long-range fragment interaction, and a region of no interaction between the two fragments. Interestingly, the range of H_{2}/CH_{3}^{+} distances spanned by each of the regions is roughly independent of the number or location of the deuterium atoms. Interestingly, the range of H_{2}/CH_{3}^{+} distances spanned by each of the regions is roughly independent of the number or location of the deuterium atoms.

aO. Asvany, S. Schlemmer, D. Gerlich, Astrophys. J. 617, 685 (2004).

bJ. B. Anderson, J. Chem. Phys. 63, 1499 (1975)

cA. B. McCoy, A. Brown, B. J. Braams, X. Huang, Z. Jin, J. M. Bowman, J. Phys. Chem. A 108, 4991 (2004)

dC. E. Hinkle, A. B. McCoy, J. Phys. Chem. A 112, 2058 (2008)

eC. E. Hinkle, A. B. McCoy, J. Phys. Chem. A 113, 4587 (2009)

fC. E. Hinkle, A. S. Petit, A. B. McCoy, in preparation

gH. S. Lee, J. M. Herbert, A. B. McCoy, J. Chem. Phys. 110, 5481 (1999)